# matrix,arrows,decorations.pathmorphing

### v2

@@ 1. Introduction
This result provides, to our knowledge, the first example of a finitely generated, infinite, residually finite, complete group whose profinite completion is also complete.
Theorem 1.3 can also be rephrased as a real anabelian property for the moduli spaces ${\mathcal{M}}_{0,n}$ of $n$-pointed, genus zero curves. Let $({\mathcal{M}}_{0,n})_{\mathbb{R}}:={\mathcal{M}}_{0,n}\times\operatorname{Spec}\mathbb{R}$, let $\pi_{1}^{\mathrm{et}}(({\mathcal{M}}_{0,n})_{\mathbb{R}})$ be its étale fundamental group for some choice of geometric base point and let $\Sigma_{n}$ be the symmetric group on $n$ letters. We then have:
@@
$\operatorname{Aut}(({\mathcal{M}}_{0,n})_{\mathbb{R}})\cong\operatorname{Out}(\pi_{1}^{\mathrm{et}}(({\mathcal{M}}_{0,n})_{\mathbb{R}}))\cong\Sigma_{n}.$
Another interesting application of Theorem 1.1 is the following. Since $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$ identifies with a subgroup of $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))$ and $G_{\mathbb{Q}}$ embeds in $\operatorname{Out}^{\mathbb{I}}({\check{\Gamma}}(S))$, we have (cf. (ii) of Proposition 4 in [16], for the genus 00 case):
Corollary 1.5.
@@
A few words about the proof of Theorem 1.1. The isomorphism $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}({\check{C}}_{P}(S))$ is a refinement of Theorem 8.1 in [6]. Here, we are able to show that there is indeed a coherent and symmetric way to define an orientation on the procongruence pants complex ${\check{C}}_{P}(S)$ (cf. Section 2.2). The proof of the identity $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ is then based on a further improvement of this isomorphism. An element $f$ of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ acts on the vertex set of ${\check{C}}_{P}(S)$, since this is naturally identified with the profinite set $\{{\hat{\mathrm{I}}}_{\sigma}|\,{\sigma}\in{\check{C}}(S)_{d(S)-1}\}$. Theorem 2.15 then states that $f$ is induced by an inner automorphism of ${\check{\Gamma}}^{\pm}(S)$ as soon as this action sends the vertex set of an edge of ${\check{C}}_{P}(S)$ to the vertex set of another edge. In order to prove the identity $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$, we then have to show that the elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\subset\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ have this property. By an induction argument, it is actually enough to show that this is true for $g(S)=0$.
@@
This is accomplished by considering the action of antiholomorphic involutions of $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ on the procongruence curve complex ${\check{C}}(S)$. By Lemma 4.2, for $g(S)=0$, there is only one ${\check{\Gamma}}^{\pm}(S)$-conjugacy class of antiholomorphic involutions in $\mathrm{P}{\check{\Gamma}}(S)$. We can then assume that a given automorphism of $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ fixes this involution and then preserves its fixed point locus in ${\check{C}}(S)$. By Lemma 4.3 and Proposition 3.6, this locus is finite and consists of isotopy classes of simple closed curves on $S$ which have between them geometric intersection either 00 or $2$. Since pairs of curves with geometric intersection $2$ correspond to edges of the pants complex, we can apply Theorem 2.15 and conclude that the given automorphism is induced by an inner automorphism of ${\check{\Gamma}}^{\pm}(S)$.
@@
Acknowledgements. We thank Javier Aramayona and Dan Margalit for some useful comments on a preliminary version of this manuscript which helped us to spot an error in the proof and improve the main result of the paper.
@@ 2.1. Augmented Teichmüller spaces and procongruence moduli stacks of curves
Let ${\mathcal{T}}(S)$ be the Teichmüller space associated to the surface $S$ and $\overline{{\mathcal{T}}}(S)$ be the augmented Teichmüller space (in [6], we called it the Bers bordification of ${\mathcal{T}}(S)$). The latter can be defined as the completion of the Teichmüller space ${\mathcal{T}}(S)$ with respect to the Weil-Petersson metric (cf. Theorem 4 in [26]). The augmented Teichmüller space $\overline{{\mathcal{T}}}(S)$ is a partial ${\Gamma}^{\pm}(S)$-equivariant compactification of ${\mathcal{T}}(S)$ such that the quotient $\overline{{\mathcal{T}}}(S)/{\Gamma}(S)$ is isomorphic to the coarse moduli space $\overline{M}(S)$ of the the DM compactification ¯${\overline{\mathcal{M}}}(S)$ of ${\mathcal{M}}(S)$.
@@
Let us denote by ~$\widetilde{{\mathcal{F}}}(S)$ the $1$-dimensional stratum of the boundary $\partial\overline{{\mathcal{T}}}(S)$. Then, each irreducible component of ~$\widetilde{{\mathcal{F}}}(S)$ is isomorphic to the cuspidal bordification ¯$\overline{{\mathbb{H}}}={\mathbb{H}}\cup{\mathbb{P}}^{1}(\mathbb{Q})$ of the hyperbolic plane ${\mathbb{H}}$ and $C_{P}(S)$ identifies with the $1$-skeleton of a ${\Gamma}(S)$-equivariant triangulation of ~$\widetilde{{\mathcal{F}}}(S)$.
@@ 2.2. Automorphisms of the procongruence pants complex and orientations
For $d(S)=1$, the pants complex $C_{P}(S)$ coincides with the Farey graph $\rm F$. This is the $1$-skeleton of a $2$-dimensional simplicial complex $\Delta$, whose geometric realization $|\Delta|$ identifies with a tessellation of the cuspidal bordification $\overline{{\mathbb{H}}}$ of the hyperbolic plane. The homeomorphism $|\Delta|\cong\overline{{\mathbb{H}}}$ becomes a conformal isomorphism when $|\Delta|$ is given the piecewise equilateral flat structure. The orientation of the Farey graph $\rm F$ (and so of $C_{P}(S)$ for $d(S)=1$) is then simply the orientation of $\Delta$ associated to the complex structure of $\overline{{\mathbb{H}}}$.
@@
The procongruence pants complex ${\check{C}}_{P}(S)$, for $d(S)>1$ is the union of the profinite set of profinite Farey graphs $\{{\widehat{\operatorname{F}}}_{\sigma}\}_{{\sigma}\in{\check{C}}(S)_{d(S)-2}}$, each naturally associated to a $(d(S)-2)$-simplex of the procongruence curve complex ${\check{C}}(S)$ (cf. Definition 6.3 in [6]), and we give each of them the orientation defined above. However, it is not clear that the automorphisms of ${\check{C}}_{P}(S)$ act in synchrony on the orientations of its profinite Farey subgraphs.
Let us denote by $O(S)$ the finite set of the topological types of $(d(S)-1)$-multicurves on $S$. To remedy the above issue, in [6], we associated a character $\operatorname{Aut}({\check{C}}_{P}(S))\to\{\pm 1\}$ to each ${\sigma}\in O(S)$ in the following way. The tautological action of $\operatorname{Aut}({\check{C}}_{P}(S))$ on ${\check{C}}_{P}(S)$ preserves profinite Farey subgraphs and ${\check{\Gamma}}(S)$-orbits of profinite Farey subgraphs. We assigned to an automorphism $\phi\in\operatorname{Aut}({\check{C}}_{P}(S))$ the plus or minus sign according to whether $\phi$ sends or not the fixed orientation of ${\widehat{\operatorname{F}}}_{\sigma}$ to the orientation of $\phi({\widehat{\operatorname{F}}}_{\sigma})$. We then proved that there is an exact sequence, for $d(S)>1$ and $S\neq S_{1,2}$ (cf. Theorem 8.1 in [6]):
(2) $1\to\operatorname{Inn}({\check{\Gamma}}(S))\to\operatorname{Aut}({\check{C}}_{P}(S))\to\prod_{O(S)}\{\pm 1\}.$
For $S=S_{1,2}$, the group $\operatorname{Aut}({\check{C}}_{P}(S_{1,2}))$ must be replaced with the subgroup of those automorphisms preserving the set of separating curves.
The first result of this section is an improved version of Theorem 8.1 in [6]. We will show that all the above characters are in fact synchronized and that the case $S\neq S_{1,2}$ is not exceptional, so that we have:
Proposition 2.1.
@@
$1\to\operatorname{Inn}({\check{\Gamma}}(S))\to\operatorname{Aut}({\check{C}}_{P}(S))\to\{\pm 1\}\to 1$
2.3. The case $S=S_{0,5}$
There is only one topological type of 00-simplices in $C(S_{0,5})$. Hence, the exact sequence (2) takes the simple form:
@@ 2.3. The case $S=S_{0,5}$
That this sequence is also right exact follows considering the action on ${\check{C}}_{P}(S_{0,5})$ of an inner automorphism $\operatorname{inn}f$ for $f\in{\Gamma}^{\pm}(S_{0,5})\smallsetminus{\Gamma}(S_{0,5})$. This proves Proposition 2.1 for $S=S_{0,5}$. Let us make some additional remarks which will be useful for the case $S=S_{1,2}$.
Let $\partial\overline{{\mathbb{M}}}(S_{0,5})$ be the DM boundary of the inverse limit $\overline{{\mathbb{M}}}(S_{0,5})$. As explained in Secion 2.1, the profinite pants complex ${\check{C}}_{P}(S_{0,5})$ identifies with the $1$-skeleton of a triangulation of $\partial\overline{{\mathbb{M}}}(S_{0,5})$. Let $\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$ be the group of automorphism which restrict to a conformal or an anticonformal map on each irreducible component and $\operatorname{Aut}^{+}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$ its subgroup consisting of conformal automorphisms.
@@
This follows from Lemma 8.8 in [6]. ∎
@@
The irreducible components of $\partial\overline{{\mathbb{M}}}(S_{0,5})$ are parameterized by the 00-simplices in ${\check{C}}(S_{0,5})$. Let us denote by $\partial\overline{{\mathbb{M}}}(S_{0,5})_{\sigma}$ the irreducible component associated to the 00-simplex ${\sigma}\in{\check{C}}(S_{0,5})_{0}$. Let us then define the character $\chi_{\sigma}\colon\thinspace\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))\to\{\pm 1\}$ which takes the value $+1$ on $f\in\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$ if $f$ sends the standard orientation of $\partial\overline{{\mathbb{M}}}(S_{0,5})_{\sigma}$ to the standard orientation of $f(\partial\overline{{\mathbb{M}}}(S_{0,5})_{\sigma})$ and $-1$ otherwise. There is an exact sequence:
@@
We can now argue exactly as in Section 8.7 of [6] and conclude that the representation $\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))\to\prod_{{\sigma}\in{\check{C}}(S_{0,5})_{0}}\{\pm 1\}$ is constant on the ${\check{\Gamma}}(S_{0,5})$-orbit of 00-simplices of ${\check{C}}(S_{0,5})$. Since there is only one such orbit, the first statement of the lemma follows. The second statement then follows from Lemma 2.2. ∎
@@
By Proposition 8.3 in [6], there is a natural monomorphism $\operatorname{Aut}({\check{C}}_{P}(S_{0,5}))\hookrightarrow\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$. The conclusion then follows from Lemma 2.3 and the fact that ${\check{\Gamma}}^{\pm}(S_{0,5})$ identifies with a subgroup of $\operatorname{Aut}({\check{C}}_{P}(S_{0,5}))$. ∎
2.4. Proof of Proposition 2.1 for $S=S_{1,2}$
For $[C,P_{1},P_{2}]\in{\mathcal{M}}_{1,[2]}$, there is a unique elliptic involution ${\upsilon}$ on $C$ and, if we denote by $C_{\upsilon}$ the (genus 00) quotient of $C$ by this involution and by $B_{\upsilon}$ the branch locus of the orbit map $C\to C_{\upsilon}$, then, the assignment $[C,P_{1},P_{2}]\mapsto[C_{\upsilon},Q,B_{\upsilon}]$, where $Q$ is the image of the pair of points $P_{1},P_{2}$ in $C_{\upsilon}$, defines a morphism of DM stacks ${\mathcal{M}}_{1,[2]}\to{\mathcal{M}}_{0,1[4]}$, where we denote by ${\mathcal{M}}_{0,1[4]}$ the moduli stack of genus 00 projective smooth curves with $5$ labeled points, one of which is singled out and the others are left unordered. The morphism ${\mathcal{M}}_{1,[2]}\to{\mathcal{M}}_{0,1[4]}$ is a $\mathbb{Z}/2$-gerbe which is split, since the composition ${\mathcal{M}}_{1,2}\to{\mathcal{M}}_{1,[2]}\to{\mathcal{M}}_{0,1[4]}$ is an isomorphism of DM stacks.

### v3

@@ 1. Introduction
Remark 1.8.
Theorem 1.7 should be compared with [18, Corollary C] which states that the group $\operatorname{Out}(\mathrm{P}{\widehat{\Gamma}}(S))$ is isomorphic to the direct product of the profinite Grothendieck-Teichmüller group with the group of permutations of the punctures (cf. also [32], for the case of profinite braid groups).
A group $G$ is complete if its center is trivial and all its automorphisms are inner, so that there is a natural isomorphism $G\cong\operatorname{Aut}(G)$. Theorem 1.7 then provides, to our knowledge, the first example of a finitely generated, infinite, residually finite, complete group whose profinite completion is also complete.
Let us recall that the étale fundamental group of an algebraic variety $X$ over $\operatorname{Spec}\mathbb{R}$ fits into the short exact sequence:
(4) $1\to\pi_{1}^{\mathrm{et}}(X_{\mathbb{C}})\to\pi_{1}^{\mathrm{et}}(X)\to G_{\mathbb{R}}\to 1,$
where we let $X_{\mathbb{C}}:=X\times\operatorname{Spec}\mathbb{C}$ and omit base points, while $G_{\mathbb{R}}\cong\mathbb{Z}/2$ is the absolute Galois group of the reals, generated by complex conjugation.
The short exact sequence (4) determines a representation $\rho_{\mathbb{R}}\colon\thinspace G_{\mathbb{R}}\to\operatorname{Out}(\pi_{1}^{\mathrm{et}}(X_{\mathbb{C}}))$ and we let $\operatorname{Out}_{G_{\mathbb{R}}}(\pi_{1}^{\mathrm{et}}(X_{\mathbb{C}}))$ be the centralizer of the image of $\rho_{\mathbb{R}}$ in $\operatorname{Out}(\pi_{1}^{\mathrm{et}}(X_{\mathbb{C}}))$. Let also $\operatorname{Aut}_{\mathbb{R}}(X)$ be the group of automorphisms of $X$ defined over $\operatorname{Spec}\mathbb{R}$. We say that $X$ satisfies the Aut-form of Grothendieck real anabelian conjecture if there is a natural isomorphism:
$\operatorname{Aut}_{\mathbb{R}}(X)\cong\operatorname{Out}_{G_{\mathbb{R}}}(\pi_{1}^{\mathrm{et}}(X_{\mathbb{C}})).$
Let us assume that the center of the group $\pi_{1}^{\mathrm{et}}(X_{\mathbb{C}})$ is trivial and that the short exact sequence (4) identifies the latter group with a characteristic subgroup of $\pi_{1}^{\mathrm{et}}(X)$. From [36, Corollary 1.5.7 and Lemma 1.6.2 (or, better, its proof)], it then follows that there is a natural isomorphism:
$\operatorname{Out}_{G_{\mathbb{R}}}(\pi_{1}^{\mathrm{et}}(X_{\mathbb{C}}))\cong\operatorname{Out}(\pi_{1}^{\mathrm{et}}(X)).$
Thus, in this case, the Aut-form of Grothendieck real anabelian conjecture can also be formulated, in an absolute form, by saying that there is a natural isomorphism:
$\operatorname{Aut}_{\mathbb{R}}(X)\cong\operatorname{Out}(\pi_{1}^{\mathrm{et}}(X)).$
Remark 1.9.
For more details on the Grothendieck anabelian conjectures and results in this direction, we refer to the survey [34], treating the case of curves, and to [17, 35] for the case of configuration spaces on curves.
Theorem 1.7 can then be rephrased as a real anabelian property for the moduli spaces ${\mathcal{M}}_{0,n}$ of $n$-pointed, genus zero curves. Let $({\mathcal{M}}_{0,n})_{\mathbb{R}}:={\mathcal{M}}_{0,n}\times\operatorname{Spec}\mathbb{R}$ and let $\pi_{1}^{\mathrm{et}}(({\mathcal{M}}_{0,n})_{\mathbb{R}})$ be its étale fundamental group for some choice of geometric base point. Let us denote by $\Sigma_{n}$ the symmetric group on $n$ letters. We have:
@@
$\operatorname{Aut}_{\mathbb{R}}(({\mathcal{M}}_{0,n})_{\mathbb{R}})\cong\operatorname{Out}(\pi_{1}^{\mathrm{et}}(({\mathcal{M}}_{0,n})_{\mathbb{R}}))\cong\Sigma_{n}.$
Remark 1.11.
The rational version of the absolute anabelian property for ${\mathcal{M}}_{0,n}$ was proved by Ihara and Nakamura in [20] (cf. Corollary C ibid.).
@@
A few words about the proof of Theorem 1.4. As we already observed above, the isomorphism $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}({\check{C}}_{P}(S))$ is a substantial refinement of [9, Theorem 8.1]. It is obtained by showing that there is indeed a coherent and symmetric way to define an orientation on the procongruence pants complex ${\check{C}}_{P}(S)$ (cf. Section 2.2). The proof of the identity $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ is then based on a further improvement of the above isomorphism. An element $f$ of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ acts on the vertex set of ${\check{C}}_{P}(S)$, since this is naturally identified with the profinite set $\{{\hat{\mathrm{I}}}_{\sigma}|\,{\sigma}\in{\check{C}}(S)_{d(S)-1}\}$. Theorem 2.17 states that $f$ is induced by an inner automorphism of ${\check{\Gamma}}^{\pm}(S)$ as soon as this action sends the vertex set of an edge of ${\check{C}}_{P}(S)$ to the vertex set of another edge. Hence, in order to prove the identities of items (ii) and (iii) of Theorem 1.4, we have to show that the elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ induced by the restriction of elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ have this property. By an induction argument, we are able to reduce to the case $g(S)=0$.
@@
We then proceed by considering the action of antiholomorphic involutions of $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ on the procongruence curve complex ${\check{C}}(S)$. By Lemma 4.5, for $g(S)=0$, there is only one ${\check{\Gamma}}^{\pm}(S)$-conjugacy class of antiholomorphic involutions in $\mathrm{P}{\check{\Gamma}}(S)$. Thus, after composing with an inner automorphism of ${\check{\Gamma}}^{\pm}(S)$, we can assume that a given automorphism of $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ fixes an antiholomorphic involution and so preserves its fixed-point locus in ${\check{C}}(S)$.
By Lemma 4.6 and Proposition 3.9, such a fixed-point locus is finite and consists of isotopy classes of simple closed curves on $S$ which have between them geometric intersection either 0 or $2$. Since pairs of curves with geometric intersection $2$ correspond to edges of the pants complex, we can apply Theorem 2.17 and conclude that the given automorphism is in fact induced by an inner automorphism of ${\check{\Gamma}}^{\pm}(S)$.
@@
Acknowledgements. We thank Javier Aramayona and Dan Margalit for some useful comments on a preliminary version of this manuscript which helped us to spot an error in the proof and improve the main result of the paper. We thank the referees for their careful reading and for their thorough comments, which led us to add many important details and improve the presentation.
@@ 2. Two preliminary results
The following two sections will be devoted to improving one of the main results of [9] (cf. [9, Theorem 8.1] and Theorem 2.2). For the convenience of the reader, we will briefly recall some notations and basic constructions from [9] and refer to that paper for more details.
@@ 2.1. Augmented Teichmüller spaces and procongruence moduli stacks of curves
Let ${\mathcal{T}}(S)$ be the Teichmüller space associated to the surface $S$ and $\overline{{\mathcal{T}}}(S)$ be the augmented Teichmüller space (in [9, Section 6.2], we called it the Bers bordification of ${\mathcal{T}}(S)$). The latter can be defined as the completion of the Teichmüller space ${\mathcal{T}}(S)$ with respect to the Weil–Petersson metric (cf. [42, Theorem 4]).
The augmented Teichmüller space $\overline{{\mathcal{T}}}(S)$ is a partial ${\Gamma}^{\pm}(S)$-equivariant compactification of ${\mathcal{T}}(S)$ such that the quotient $\overline{{\mathcal{T}}}(S)/{\Gamma}(S)$ is isomorphic to the coarse moduli space $\overline{M}(S)$ of the DM compactification ${\overline{\mathcal{M}}}(S)$ of ${\mathcal{M}}(S)$.
@@
Let us denote by $\widetilde{{\mathcal{F}}}(S)$ the $1$-dimensional stratum of the boundary $\partial\overline{{\mathcal{T}}}(S)$. Then, each irreducible component of $\widetilde{{\mathcal{F}}}(S)$ is isomorphic to the cuspidal bordification $\overline{{\mathbb{H}}}={\mathbb{H}}\cup{\mathbb{P}}^{1}(\mathbb{Q})$ of the hyperbolic plane $\mathbb{H}$} and $C_{P}(S)$ identifies with the $1$-skeleton of a ${\Gamma}(S)$-equivariant triangulation of $\widetilde{{\mathcal{F}}}(S)$.
@@ 2.2. Automorphisms of the procongruence pants complex and orientations
For $d(S)=1$, the pants complex $C_{P}(S)$ coincides with the Farey graph $\rm F$. This is the $1$-skeleton of a $2$-dimensional simplicial complex $\Delta$, whose geometric realization $|\Delta|$ is identified with a tessellation of the cuspidal bordification $\overline{{\mathbb{H}}}$ of the hyperbolic plane. The homeomorphism $|\Delta|\cong\overline{{\mathbb{H}}}$ becomes a conformal isomorphism when $|\Delta|$ is given the piecewise equilateral flat structure. The orientation of the Farey graph $\rm F$ (and hence of $C_{P}(S)$ for $d(S)=1$) is then simply the orientation of $\Delta$ associated to the complex structure of $\overline{{\mathbb{H}}}$.
@@
The procongruence pants complex ${\check{C}}_{P}(S)$, for $d(S)>1$ is the union of the profinite set of profinite Farey graphs $\{{\widehat{\operatorname{F}}}_{\sigma}\}_{{\sigma}\in{\check{C}}(S)_{d(S)-2}}$, each naturally associated to a (d(S)2)(d(S)-2)-simplex of the procongruence curve complex ${\check{C}}(S)$ (cf. [9, Definition 6.3]), and we give each of them the orientation defined above. However, it is not clear whether the automorphisms of ${\check{C}}_{P}(S)$ act in synchrony on the orientations of its profinite Farey subgraphs.
Let us denote by $O(S)$ the finite set of the topological types of (d(S)1)(d(S)-1)-multicurves on $S$. To remedy the above issue, in [9], we associated a character $\chi_{\sigma}\colon\thinspace\operatorname{Aut}({\check{C}}_{P}(S))\to\{\pm 1\}$ to each ${\sigma}\in O(S)$ in the following way. The tautological action of $\operatorname{Aut}({\check{C}}_{P}(S))$ on ${\check{C}}_{P}(S)$ preserves profinite Farey subgraphs and ${\check{\Gamma}}(S)$-orbits of profinite Farey subgraphs. We assigned to an automorphism $\phi\in\operatorname{Aut}({\check{C}}_{P}(S))$ the plus or minus sign according to whether $\phi$ sends or not the fixed orientation of ${\widehat{\operatorname{F}}}_{\sigma}$ to the orientation of $\phi({\widehat{\operatorname{F}}}_{\sigma})$. We then proved that there is an exact sequence, for $d(S)>1$ and $S\neq S_{1,2}$ (cf. [9, Theorem 8.1]):
(5) $1\to\operatorname{Inn}({\check{\Gamma}}(S))\to\operatorname{Aut}({\check{C}}_{P}(S))\to\prod_{O(S)}\{\pm 1\},$
where, for $S=S_{1,2}$, the group $\operatorname{Aut}({\check{C}}_{P}(S_{1,2}))$ has to be replaced with the subgroup of those automorphisms preserving the set of separating curves.
Remark 2.1.
It is easy to check that all characters $\chi_{\sigma}\colon\thinspace\operatorname{Aut}({\check{C}}_{P}(S))\to\{\pm 1\}$, for ${\sigma}\in O(S)$, are nontrivial. For this, let us observe that the action of an element $f\in{\Gamma}^{\pm}(S)\smallsetminus{\Gamma}(S)$ on the augmented Teichmüller space $\overline{{\mathcal{T}}}(S)$ inverts the orientation of every open stratum of this space. This implies that, for all ${\sigma}\in O(S)$, we have $\chi_{\sigma}(\operatorname{inn}f)=-1$.
The first result of this section is an improved version of [9, Theorem 8.1]. We will show that all the above characters are in fact synchronized and that the case $S=S_{1,2}$ is not exceptional, so that we have:
Theorem 2.2.
@@
$1\to\operatorname{Inn}({\check{\Gamma}}(S))\to\operatorname{Aut}({\check{C}}_{P}(S))\to\{\pm 1\}\to 1.$
2.3. Proof of Theorem 2.2 for $S=S_{0,5}$
There is only one topological type of 0-simplices in $C(S_{0,5})$. Hence, the exact sequence (5) takes the simple form:
@@
From Remark 2.1, it then immediately follows that the above sequence is also right exact, which proves Theorem 2.2 for $S=S_{0,5}$.
2.4. Proof of Theorem 2.2 for $S=S_{1,2}$
To deal with this case, we first need to analyze in more detail the case $S=S_{0,5}$. Let $\partial\overline{{\mathbb{M}}}(S_{0,5})$ be the DM boundary of the inverse limit $\overline{{\mathbb{M}}}(S_{0,5})$. As explained in Section 2.1, the profinite pants complex ${\check{C}}_{P}(S_{0,5})$ identifies with the $1$-skeleton of a triangulation of $\partial\overline{{\mathbb{M}}}(S_{0,5})$. Let $\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$ be the group of automorphisms which restrict to a conformal or an anticonformal map on each irreducible component and $\operatorname{Aut}^{+}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$ its subgroup consisting of conformal automorphisms.
@@
For simplicity, put $S=S_{0,5}$ and, for levels ${\Gamma}^{\lambda}\subseteq{\Gamma}^{\mu}\subseteq{\Gamma}(S)$, let us denote by $\hat{\pi}_{\lambda}\colon\thinspace\partial\overline{{\mathbb{M}}}(S)\to\partial\overline{{\mathcal{M}}}(S)^{\lambda}$ and $\pi_{{\lambda}\mu}\colon\thinspace\partial\overline{{\mathcal{M}}}(S)^{\lambda}\to\partial\overline{{\mathcal{M}}}(S)^{\mu}$ the natural projections. Then, a conformal automorphism $\phi$ of $\partial\overline{{\mathbb{M}}}(S)$ determines, and is determined by, the directed inverse system of conformal maps constructed as follows. For every ${\lambda}\in{\Lambda}$, there is a $\mu\in{\Lambda}$ such that the composition $\phi_{\lambda}:=\hat{\pi}_{\lambda}\circ\phi\colon\thinspace\partial\overline{{\mathbb{M}}}(S)\to\partial\overline{{\mathcal{M}}}(S)^{\lambda}$ factors through a conformal map $\phi_{\mu{\lambda}}\colon\thinspace\overline{{\mathcal{M}}}(S)^{\mu}\to\overline{{\mathcal{M}}}(S)^{\lambda}$. It is clear that, if $\phi_{\mu^{\prime}{\lambda}^{\prime}}\colon\thinspace\overline{{\mathcal{M}}}(S)^{\mu^{\prime}}\to\overline{{\mathcal{M}}}(S)^{{\lambda}^{\prime}}$ is another such map with ${\Gamma}^{{\lambda}^{\prime}}\subseteq{\Gamma}^{\lambda}$ and ${\Gamma}^{\mu^{\prime}}\subseteq{\Gamma}^{\mu}$, we have $\pi_{{\lambda}^{\prime}{\lambda}}\circ\phi_{\mu^{\prime}{\lambda}^{\prime}}=\phi_{\mu{\lambda}}\circ\pi_{\mu^{\prime}\mu}$. Therefore, the set of conformal maps $\{\phi_{\mu{\lambda}}\}_{{\lambda},\mu\in{\Lambda}}$ is a directed inverse system with inverse limit the map $\phi$.
Let us now observe that, for $S=S_{0,5}$, the DM boundary $\partial\overline{{\mathcal{M}}}(S)^{\lambda}$ of the level structure $\overline{{\mathcal{M}}}(S)^{\lambda}$ coincides with the Fulton curve ${\mathcal{F}}^{\lambda}\subset\overline{{\mathcal{M}}}(S)^{\lambda}$ (cf. [9, Definition 6.1]). The conclusion then follows from [9, Lemma 8.8]. ∎
@@
The irreducible components of $\partial\overline{{\mathbb{M}}}(S_{0,5})$ are parameterized by the 0-simplices in ${\check{C}}(S_{0,5})$. Let us denote by $\partial\overline{{\mathbb{M}}}(S_{0,5})_{\sigma}$ the irreducible component associated to the 0-simplex ${\sigma}\in{\check{C}}(S_{0,5})_{0}$. Let us then define the character $\chi_{\sigma}\colon\thinspace\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))\to\{\pm 1\}$ which takes the value +1+1 on $f\in\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$ if $f$ sends the standard orientation of $\partial\overline{{\mathbb{M}}}(S_{0,5})_{\sigma}$ to the standard orientation of $f(\partial\overline{{\mathbb{M}}}(S_{0,5})_{\sigma})$ and 1-1 otherwise. There is an exact sequence:
@@
We can now argue exactly as in [9, Section 8.7] and conclude that the representation $\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))\to\prod_{{\sigma}\in{\check{C}}(S_{0,5})_{0}}\{\pm 1\}$ is constant on the ${\check{\Gamma}}(S_{0,5})$-orbit of 0-simplices of ${\check{C}}(S_{0,5})$. Since there is only one such orbit, the first statement of the lemma follows. The second statement then follows from Lemma 2.3. ∎
@@
As we remarked above, ${\check{C}}_{P}(S_{0,5})$ identifies with the $1$-skeleton of a triangulation of $\partial\overline{{\mathbb{M}}}(S_{0,5})$. By [9, Theorem 6.7 and Proposition 8.3], there is then a natural monomorphism $\operatorname{Aut}({\check{C}}_{P}(S_{0,5}))\hookrightarrow\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$. Thus, the conclusion follows from Lemma 2.4 and the fact that ${\check{\Gamma}}^{\pm}(S_{0,5})$ identifies with a subgroup of $\operatorname{Aut}({\check{C}}_{P}(S_{0,5}))$. ∎
For $C,P_{1},P_{2}]\in{\mathcal{M}}_{1,[2]}$, there is a unique elliptic involution $\upsilon$} on $C$ and, if we denote by $C_{\upsilon}$ the (genus 0) quotient of $C$ by this involution and by $B_{\upsilon}$ the branch locus of the orbit map $C\to C_{\upsilon}$, then, the assignment $C,P_{1},P_{2}]\mapsto[C_{\upsilon},Q,B_{\upsilon}]$, where $Q$ is the image of the pair of points $P_{1},P_{2}$ in $C_{\upsilon}$, defines a morphism of DM stacks ${\mathcal{M}}_{1,[2]}\to{\mathcal{M}}_{0,1[4]}$, where we denote by ${\mathcal{M}}_{0,1[4]}$ the moduli stack of genus 0 projective smooth curves with $5$ labeled points, one of which is singled out and the others are left unordered. The morphism ${\mathcal{M}}_{1,[2]}\to{\mathcal{M}}_{0,1[4]}$ is a $\mathbb{Z}/2$-gerbe which is split, since the composition ${\mathcal{M}}_{1,2}\to{\mathcal{M}}_{1,[2]}\to{\mathcal{M}}_{0,1[4]}$ is an isomorphism of DM stacks. This can be checked observing that the morphism ${\mathcal{M}}_{1,2}\to{\mathcal{M}}_{0,1[4]}$ induces an isomorphism both on the respective coarse moduli spaces and on the isotropy groups of points.